# Robust, fast, and efficient formation of stable tetratomic molecules from ultracold atoms via generalized stimulated Raman exact passage

### v3

@@ III.2 Result and Discussion
On the other hand, Figure. 8(b) presents a contour plot of the transfer efficiency as a function of $\delta\tilde{\Omega}_{2}$ and $\delta\tilde{\Omega}_{3}$, with Condition (18) maintained; in this case, $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ will passively vary with the other two ones. It is clear that the transfer efficiency can keep at a high level, even when the deviation of the Rabi frequency deviations within $|\delta\tilde{\Omega}_{2}/\tilde{\Omega}_{2}|\!=\!|\delta\tilde{\Omega}_{3}/\tilde{\Omega}_{3}|\!=\!10\%$. This result further demonstrates the necessity of Condition (18).
@@ IV Conclusions and Outlooks
To conclude, we present a theoretical formalism for the formation of stable ultracold tetratomic molecules. The process begins with a mixture of ultracold atoms: first, generalized nonlinear STIREP is employed to coherently assemble atoms into tetratomic molecules; subsequently, to ensure the resulting molecules exhibit sufficient stability, we propose that the C-STIREP technique be used to transfer them to the stable ground state. Numerical analyses show that the proposed two-step strategy enables the production of stable ultracold molecules with nearly ideal efficiency, while featuring excellent accelerability and high robustness.
The main advantages of our work are organized as follows: (i) It paves a potential pathway for forming stable ultracold tetratomic molecules, and in principle, has the potential to be extended to preparing larger polyatomic molecular complexes Dou et al. (2013); (ii) Unlike traditional methods, the adiabatic condition is released in our approach, thus enabling fast speed and high efficiency. Additionally, the transient populations of short-lived states can be effectively suppressed without relying on the dark state; (iii) Although the protocol follows the spirit of STA, it does not introduce extra couplings, and the applied pulses pose no actual implementation difficulties; (iv) Numerous exact solutions support the achievement of predefined goals, and the above-discussed examples (7) and (20) are not exhaustive Chen and Muga (2012); Kang et al. (2016); Baksic et al. (2016); Wu et al. (2019).
Finally, we point out that the tetratomic molecule preparation process suggested in this study aligns closely with the standard procedure for diatomic molecule, with both following the similar route of “first creating unstable molecular samples by assembling pre-cooled atom, and then converting them to sufficiently stable ground state”. Inspired by the studies in Refs. Mackie and Phou (2010); Price and Yelin (2019), we will, in the future, commit to exploring a coherent strategy that starts directly from ultracold atoms and enables the robust, fast and efficient production of stable UPMs. This problem may be worth mentioning.
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grants No. 12475026, No. 12075193 and No. 12365004) and the Doctoral Research Startup Fund of Northwest Normal University (Grant No. 6014/202503101301).

### v4

@@ III.2 Result and Discussion
Figure. 9(b) presents a contour plot of the transfer efficiency as a function of $\delta\tilde{\Omega}_{2}$ and $\delta\tilde{\Omega}_{3}$, with Condition (20) maintained; in this case, $\tilde{\Omega}_{1}$ and $\tilde{\Omega}_{4}$ will passively vary with the other two. It is clear that the transfer efficiency can keep at a high level, when the deviation of the Rabi frequency deviations is within $|\delta\tilde{\Omega}_{2}/\tilde{\Omega}_{2}|\!=\!|\delta\tilde{\Omega}_{3}/\tilde{\Omega}_{3}|\!=\!10\%$. This result further demonstrates the necessity of Condition (20).
@@ IV Conclusions and Outlooks
To conclude, we present a theoretical formalism for the formation of stable ultracold tetratomic molecules. The process begins with a mixture of ultracold atoms: first, generalized nonlinear STIREP is employed to coherently assemble atoms into tetratomic molecules; subsequently, to ensure the resulting molecules exhibit sufficient stability, we propose that the C-STIREP technique be used to transfer them to the stable ground state. Numerical analyses show that the proposed two-step strategy enables the production of stable ultracold molecules with nearly ideal efficiency, while featuring excellent accelerability and high robustness. Although this is a theoretical study, we hope that with the advancement of experimental techniques, particularly in the formation of long-lived triatomic and tetratomic molecules, this two-step strategy will be translated into practical applications in the years to come Wunsch et al. (2011); Kunitski et al. (2015); Yang et al. (2022); Mukherjee and Tomza (2026).
The main advantages of our work are: (i) It paves a potential pathway for forming stable ultracold tetratomic molecules, and in principle, has the potential to be extended to preparing larger polyatomic molecular complexes Dou et al. (2013); (ii) Unlike traditional methods, the adiabatic condition is released in our approach, thus enabling fast speed and high efficiency. Additionally, the transient populations of short-lived states can be effectively suppressed without relying on the dark state; (iii) Although the protocol follows the spirit of STA, it does not introduce extra couplings, and the applied pulses pose no actual implementation difficulties; (iv) Numerous exact solutions support the achievement of predefined goals, and the above-discussed examples (9) and (22) are not exhaustive.
Finally, we point out that the tetratomic molecule preparation process suggested in this study aligns closely with the standard procedure for diatomic molecule, with both following the similar route of “first creating unstable molecular samples by assembling pre-cooled atom, and then converting them to sufficiently stable ground state”. Inspired by the studies in Refs. Mackie and Phou (2010); Price and Yelin (2019), we will, in the future, commit to exploring a coherent strategy that starts directly from ultracold atoms and enables the robust, fast and efficient production of stable UPMs.
@@ Acknowledgments
We would like to thank the anonymous referees for constructive comments that are helpful for improving the quality of the work. This work was supported by the National Natural Science Foundation of China (Grants No. 12475026, No. 12075193, and No. 12365004) and the Doctoral Research Startup Fund of Northwest Normal University (Grant No. 6014/202503101301).
Appendix A
In the Appendix, we will present the detailed derivation process from Eq. (2) to (6). According to Eq. (3), we have
$\begin{aligned} \dot{\psi}_{\mathrm{a}}&\!=\!\dot{\psi}_{\mathrm{a}}e^{i\alpha(t)}\!+\!i\dot{\alpha}(t)\psi_{\mathrm{a}}e^{i\alpha(t)},\\ \dot{\psi}_{\mathrm{m}}&\!=\!\dot{\psi}_{\mathrm{m}}e^{3i\alpha(t)}\!+\!3i\dot{\alpha}(t)\psi_{\mathrm{m}}e^{3i\alpha(t)},\\ \dot{\psi}_{\mathrm{t}}&\!=\!\dot{\psi}_{\mathrm{t}}e^{4i\alpha(t)}\!+\!4i\dot{\alpha}(t)\psi_{\mathrm{t}}e^{4i\alpha(t)}.\end{aligned}$ (24)
Substituting the above results into Eq. (2), we obtain
$\begin{aligned} i\left[\dot{\psi}_{\mathrm{a}}e^{i\alpha\left(t\right)}\!+\!i\dot{\alpha}\left(t\right)\psi_{\mathrm{a}}e^{i\alpha\left(t\right)}\right]\!&=\!-2K_{\mathrm{a}}\left(t\right)\psi_{\mathrm{a}}e^{i\alpha\left(t\right)}\!-\!3\Omega_{1}{\psi_{\mathrm{m}}e^{3i\alpha\left(t\right)}\psi^{\ast}_{\mathrm{a}}}^{2}e^{-2i\alpha\left(t\right)}\!+\!\Omega_{2}\psi_{\mathrm{t}}e^{4i\alpha\left(t\right)}{\psi^{*}_{\mathrm{m}}}e^{-3i\alpha\left(t\right)},\\ i\left[\dot{\psi}_{\mathrm{m}}e^{3i\alpha\left(t\right)}\!+\!3i\dot{\alpha}\left(t\right)\psi_{\mathrm{m}}e^{3i\alpha\left(t\right)}\right]&=\!-2K_{\mathrm{m}}\left(t\right)\psi_{\mathrm{m}}e^{3i\alpha\left(t\right)}\!-\!\left(\delta\!+\!i\gamma_{\mathrm{m}}\right)\psi_{\mathrm{m}}e^{3i\alpha\left(t\right)}\!-\!\Omega_{1}\psi_{\mathrm{a}}^{3}e^{3i\alpha\left(t\right)}\!+\!\Omega_{2}\psi_{\mathrm{t}}e^{4i\alpha\left(t\right)}{\psi^{\ast}_{\mathrm{a}}e^{-i\alpha\left(t\right)}},\\ i\left[\dot{\psi}_{\mathrm{t}}e^{4i\alpha\left(t\right)}\!+\!4i\dot{\alpha}\left(t\right)\psi_{\mathrm{t}}e^{4i\alpha\left(t\right)}\right]\!&=\!-2K_{\mathrm{t}}\left(t\right)\psi_{\mathrm{t}}e^{4i\alpha\left(t\right)}\!-\!\left(\Delta\!+\!\delta\!+\!i\gamma_{\mathrm{t}}\right)\psi_{\mathrm{t}}e^{4i\alpha\left(t\right)}\!+\!\Omega_{2}\psi_{\mathrm{m}}e^{3i\alpha\left(t\right)}\psi_{\mathrm{a}}e^{i\alpha\left(t\right)}.\end{aligned}$ (25)
It is evident that all terms on both sides of the above equation share a common factor of $e^{i\alpha(t)}$. By canceling out this common factor, the preceding equation can be simplified to:
$\begin{aligned} i\dot{\psi}_{\mathrm{a}}\!-\!\dot{\alpha}(t)\psi_{\mathrm{a}}\!&=\!-2K_{a}(t)\psi_{\mathrm{a}}\!-\!3\Omega_{1}{\psi_{\mathrm{m}}\psi^{\ast}_{\mathrm{a}}}^{2}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{*}_{\mathrm{m}}},\\ i\dot{\psi}_{\mathrm{m}}\!-\!3\dot{\alpha}(t)\psi_{\mathrm{m}}\!&=\!-2K_{\mathrm{m}}(t)\psi_{\mathrm{m}}\!-\!(\delta\!+\!i\gamma_{\mathrm{m}})\psi_{\mathrm{m}}\!-\!\Omega_{1}\psi_{\mathrm{a}}^{3}\!+\!\Omega_{2}\psi_{\mathrm{t}}{\psi^{\ast}_{\mathrm{a}}},\\ i\dot{\psi}_{\mathrm{t}}\!-\!4\dot{\alpha}(t)\psi_{\mathrm{t}}\!&=\!-2K_{\mathrm{t}}(t)\psi_{\mathrm{t}}\!-\!(\Delta\!+\!\delta\!+\!i\gamma_{\mathrm{t}})\psi_{\mathrm{t}}\!+\!\Omega_{2}\psi_{\mathrm{m}}\psi_{\mathrm{a}}.\end{aligned}$ (26)
After moving all the like terms to one side, we can arrive at Eq. (4) presented in Section. II.1. Finally, by setting $\dot{\alpha}=2K_{a}(t)$ and performing some straightforward calculations, we end up with the condition for resonance locking (5); once this condition is satisfied, the system can be simplified to Eq. (6).
Appendix B
For any physical quantity $X$, its deviation is defined as
$\begin{aligned} \delta X=X\times\delta X/X,\end{aligned}$ (27)
where $X$ denotes its ideal value, and $\delta X/X$ represents the dimensionless relative deviation (ranging from 0.1-0.1 to $0.1$, i.e., $\pm 10\%$). After incorporating the deviation, the realistic physical quantity used for calculation is given by $X\!+\!\delta X$.